Required information e following information applies to the questions displayed below. n Co. reports the contribution margin income statement for 2017 HUDSON CO Contribution Margin Income Statement For Year Ended December 31, 2017 Sales (10,300 units at $375 each) Variable costs (10, 300 units at $300 each) Contribution margin Fixed costs Pretax income $3,862,500 3,090,000 $ 772,500 600,000 $ 172,500 1. Compute Hudson Co.s break-even point in units and. 2. Compute Hudson Co.s break-even point in sales dollars 1. Break-even point 2. Break-even point units

3.4.

5.

6.
Answer #1

1

Contribution margin ratio
Ref Particulars Amount
a Sales revenue        3,862,500
b Contribution           772,500
c=b/a *100 Contribution margin ratio 20%
Break even point
a Contribution per unit               75.00
b Fixed costs           600,000
c=b/a BEP units          8,000.00

2

Break even point
Ref Particulars Amount
a Contribution margin ratio 20%
b Fixed costs           600,000
c=b/a BEP sales        3,000,000

3

Particulars Amount
Required profit           169,000
Fixed costs           600,000
Required contribution           769,000
CMR 20%
Required sales        3,845,000

Margin of safety

Margin of safety
Ref Particulars Amount
a Current sales        3,845,000
b BE sales        3,000,000
c= a-b Margin sales           845,000
d= c/a Margin of safety 21.98%

4

Income statement
Particulars Per unit Total
Units             10,300
Sales revenue             375.00        3,862,500
Minus:
Variable cost             290.00        2,987,000
Contribution               85.00           875,500
Minus:
Fixed cost           644,000
Net operating income           231,500

Company should purchase as it increases the profit by 59000

5

Income statement
Particulars Per unit Total
Units             10,300
Sales revenue             400.00        4,120,000
Minus:
Variable cost             300.00        3,090,000
Contribution margin             100.00        1,030,000
Minus:
Fixed cost           600,000
Net operating income           430,000
Contribution margin ratio
Ref Particulars Amount
a Sales revenue        4,120,000
b Contribution margin        1,030,000
c=b/a *100 Contribution margin ratio 25%
Break even point
a Contribution per unit             100.00
b Fixed costs           600,000
c=b/a BEP units          6,000.00
Break even point
a Contribution margin ratio 25%
b Fixed costs           600,000
c=b/a BEP sales        2,400,000

6

Income statement
Particulars Per unit Total
Units             11,700
Sales revenue             375.00        4,387,500
Minus:
Variable cost             300.00        3,510,000
Contribution margin               75.00           877,500
Minus:
Fixed cost           765,000
Net operating income           112,500
Existing income           172,500
Change            -60,000

It decrease net income, should not spend for the proposed expenditure.